Exercise 15-24 (Static) Calculation of annual lease payments; residual value [LO15-2, 15-6]

Each of the four independent situations below describes a finance lease in which annual lease payments are payable at the beginning of each year. The lessee is aware of the lessor’s implicit rate of return. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Check 10 1 4 10% $50,000 $50,000 Lease term (years) Lessors rate of return Fair value of lease asset Lessors cost of lease
Answer #1
Lease Payments Residual Value Guarantee PV Of Lease Payments PV Of Residual Value Guarantee Right-of-use Asset/ Lease Liability
Situation 1 $ 14,340 $ 0 $ 50,000 $ 0 $ 50,000
Situation 2 $ 62,310 $ 0 $ 325,917 $ 0 $ 325,917
Situation 3 $ 16,617 $ 0 $ 70,450 $ 0 $ 70,450
Situation 4 $ 79,947 $ 5,000 {$50,000-$45,000} $ 444,806 $2,019 $ 446,825 {$444,806+ $2,019}
Explanation :
Situation 1
Lease Payments = $ 50,000 / 3.48685 = $ 14,340
Lease Payments = $ 14,340 (Rounded )
Present value of annuity due factor $ 1 for n= 4 , i= 10% is 3.48685
Situation 2
Lease Payments = (Fair Value - Present value of Residual Value )/ Present Value of Annuity due Factor
Present Value of Residual Value = $ 50,000 * 0.48166 = $24,083
Present value of $ 1 , n=7 , i= 11% is 0.48166
Present value of annuity due factor $ 1 for n= 7 , i= 11% is 5.23054
Lease Payments = (Fair Value - Present value of Residual Value )/ Present Value of Annuity due Factor
Lease Payments = ($350,000 - $24,083)/ 5.23054
Lease Payments = $ 62,310 (Rounded)
Present Value of Lease Payments = (Fair Value - Present value of Residual Value )
Present Value of Lease Payments = ($350,000 - $24,083)
Present Value of Lease Payments = $ 325,917
Situation 3
Lease Payments = (Fair Value - Present value of Residual Value )/ Present Value of Annuity due Factor
Present Value of Residual Value = $ 7,000 * 0.64993 = $ 4,550 (Rounded)
Present value of $ 1 , n=5 , i= 9% is 0.64993
Present value of annuity due factor $ 1 for n= 5 , i= 9% is 4.23972
Lease Payments = (Fair Value - Present value of Residual Value )/ Present Value of Annuity due Factor
Lease Payments = ($75,000 - $4,550)/ 4.23972
Lease Payments = $ 16,617 (Rounded)
Present Value of Lease Payments = (Fair Value - Present value of Residual Value )
Present Value of Lease Payments = ($75,000 - $4,550)
Present Value of Lease Payments = $ 70,450
Situation 4
Lease Payments = (Fair Value - Present value of Guaranteed Residual Value )/ Present Value of Annuity due Factor
Present Value of Guaranteed Residual Value = $ 50,000 * 0.40388 = $ 20,194 (Rounded)
Present value of $ 1 , n=8 , i= 12% is 0.40388
Present value of annuity due factor $ 1 for n= 8 , i= 12% is 5.56376
Lease Payments = (Fair Value - Present value of Guaranteed Residual Value )/ Present Value of Annuity due Factor
Lease Payments = ($465,000 - $20,194)/ 5.56376
Lease Payments = $ 79,947 (Rounded)
Present Value of Lease Payments = (Fair Value - Present value of Guaranteed Residual Value )
Present Value of Lease Payments = ($465,000 - $20,194)
Present Value of Lease Payments = $ 444,806
Present Value of Residual Value Guarantee = ($50,000 -$45,000)* 0.40388
Present Value of Residual Value Guarantee = $5,000 * 0.40388
Present Value of Residual Value Guarantee = $ 2,019 (Rounded)